Berjaya publishing publish two type of magazines, namely sports magazines and car magazines. The sports magazine (SM) is sold in large quantities to various sports centers around the country. The car magazine (CM) on the other had is a slow moving item, therefore produced on monthly basis. Due to its low volume, only a small number of inspection are carried out during production. However, the CM relies on colorful car pictures and uses international information to attract its readers. The CM therefore uses more machine hours than the SM in its production. The company currently uses traditional absorption costing system based on machine hours to account for its overhead. The following information is provided for the current accounting year. SM RM per unit CM RM per unit Direct production cost: 1) Paper 2) Printing ink 3) Machine Cost 1.50 0.16 2.90 8.94 2.30 3.90 Overhead cost 4.6 7.90 Total cost Per unit 11.30 20.90 The managing directors has recently attended a seminar on activity based costing (ABC) and he instructed you to implement ABC approach for the coming year. The following information is provided for the coming year. 1) The paper used in SM production cost RM 4 per kg but the CM paper costs only RM 2 per kg. the SM uses 400 grams of paper for each magazine, four times as much as the CM uses. 2) Printing ink costs RM 60 per liter. The SM uses one third of the printing ink of the more colorful CM. the CM uses 150ml of printing ink per magazine. 3) The CM needs twelve minutes of machine time to produce each magazine, whereas the CM needs twenty minutes per magazine. The machines cost RM12 per hour to run. 4) There are three main overheads, the data for there are: Overhead annual Cost for the coming year RM Activity driver Property costs 4,320,000 Machine hours Quality control 1,336,000 Number of inspections Production set up costs 104,000 Number of production runs Total 5,760,000 5) The SM will be inspected on 180 occasions next year, whereas the CM will be inspected just 20 times. 6) Berjaya publishing will produce its annual output of 2,000,000 SMs in eight production runs and approximately 240,000 CMs in 12producio runs. Required: a) Calculate the cost per unit for the SM and the CM using machine hours to absorb the overheads. b) Calculate the cost per unit for the SM and the CM using activity based costing principles to absorb the overheads.
Answer #1
a)Cost per unit for the SM and the CM using machine hours to absorb the overheads.
Workings: SM CM
Output(no.of magazines) 2000000 240000
Total M/c hrs. reqd.
2000000*12 min./60 min. 400000
240000*20/60= 80000
Allocation of OH:
5760000/480000*400000= 4800000
5760000/480000*80000= 960000
OH/Unit(Total allocated/output unit)
4800000/2000000= 2.4
960000/240000= 4
Cost per unit: SM RM/unit CM RM/unit
Direct production cost
1. Paper 4/1000*400= 1.6 2/1000*100= 0.2
2.Printing Ink 60/1000*50= 3 60/1000*150= 9
3.Machine cost 12/60*12= 2.4 12/60*20= 4
4.OH based on m/c hrs.(as per above Wkgs.) 2.4 4
Total cost per unit 9.4 17.2
b)Cost per unit for the SM and the CM using activity based costing principles to absorb the overheads
Workings per GIVEN data: ALLOCATION
Annual OH cost for the coming year RM Activity driver SM CM SM CM
Property costs 4320000 M/c hrs. 400000 80000 4320000/480000*400000= 3600000 4320000/480000*80000= 720000
Quality control 1336000 no.of inspecns. 180 20 1336000/200*180= 1202400 1336000/200*20= 133600
Prodn. Set-up costs 104000 no.of prodn.runs 8 12 104000/20*8= 41600 104000/20*12= 62400
Total OH cost 5760000 4844000 916000
Output(no.of magazines) 2000000 240000 2000000 240000
OH Per unit(as per ABC principles) 2.42 3.82
(Total Ohs as per ABC allocation/No.of output units)
Cost per unit (ABC) SM RM/unit CM RM/unit
Direct production cost
1. Paper 4/1000*400= 1.6 2/1000*100= 0.2
2.Printing Ink 60/1000*50= 3 60/1000*150= 9
3.Machine cost 12/60*12= 2.4 12/60*20= 4
4. OH /unit as per ABC(as per above Wkgs.) 2.42 3.82
Total cost per unit 9.42 17.02